(𝑑 +2)^2 𝑦′′ + (𝑑+2) 𝑦′ + 𝑦=0



1
Expert's answer
2021-06-13T17:19:53-0400

"\displaystyle
(t + 2)^2 y'' + (t + 2)y' + y = 0\\

y = f(\ln(t + 2)) = f(x)\\

y' = \frac{f'(\ln(t + 2))}{t + 2} \\

y'' = \frac{f''(\ln(t + 2))}{(t + 2)^2} - \frac{f'(\ln(t + 2))}{(t + 2)^2}\\

f''(\ln(t + 2)) - f'(\ln(t + 2)) + f'(\ln(t + 2)) + f(\ln(t + 2)) = 0\\

f''(\ln(t + 2)) + f(\ln(t + 2)) = 0\\

f''(x) + f(x) = 0\\


f(x) = c_1\cos{x} + c_2\sin{x}\\

f(x) = c_1\cos(\ln(t + 2)) + c_2\sin(\ln(t + 2))"